### v1

# LUPET: Incorporating Hierarchical Information Path into Multilingual ASR

Abstract
Many factors have separately shown their effectiveness on improving multilingual ASR. They include language identity (LID) and phoneme information, language-specific processing modules and cross-lingual self-supervised speech representation, etc. However, few studies work on synergistically combining them to contribute a unified solution, which still remains an open question. To this end, a novel view to incorporate hierarchical information path LUPET into multilingual ASR is proposed. The LUPET is a path encoding multiple information in different granularity from shallow to deep encoder layers. Early information in this path is beneficial for deriving later occurred information. Specifically, the input goes from LID prediction to acoustic unit discovery followed by phoneme sharing, and then dynamically routed by mixture-of-expert for final token recognition. Experiments on 10 languages of Common Voice examined the superior performance of LUPET. Importantly, LUPET significantly boosts the recognition on high-resource languages, thus mitigating the compromised phenomenon towards low-resource languages in a multilingual setting.
Index Terms: Multilingual ASR, language identity, self-supervised speech representation learning, mixture-of-expert
1 Introduction
Multilingual ASR aims to use one unified model to transcribe speech in multiple languages to their textual representations [1, 2, 3, 4]. This offers two great advantages [5] that (1) the per-language training cost and deployment effort are largely reduced; (2) recognition performance on low-resource languages can be improved a lot, benefiting from the shared language-invariant knowledge. Previous studies [6] showed that directly training a fully-shared end-to-end (E2E) model on a multilingual corpus is a simple yet effective solution (vanilla). The multilingual corpus is formed by mixing corpora of different languages, meanwhile creating a shared vocabulary set. However, due to the heterogeneous nature among languages [7, 6], the vanilla scheme often causes high-resource languages to compromise their performance towards low-resources in multilingual training, i.e., worse than their monolingual counterparts.
To mitigate the degradation on high-resource languages, recent research considers incorporating different multilingual-related information, e.g., language identity (LID) [8, 9, 10, 11] and phoneme [12, 13, 14] or language-specific architecture, e.g., mixture-of-expert (MoE) [15, 16, 17, 18, 19]. Either including the above information or adopting MoE architecture, one common feature is that they are both in a supervised training framework. Self-supervised learning (SSL) has presented numerous successful practices, like wav2vec2.0 [20], HuBERT [21] and XLSR [22], etc, on speech recognition and should not be neglected. Usually, a two-stage training scheme, i.e., pre-train and then fine-tune, is widely applied. In [23], it was found that jointly unsupervised and supervised training (JUST) on multilingual ASR can outperform the two-stage scheme and prevent forgetting knowledge learned by SSL. The SSL in JUST consists of a contrastive loss and a masked language model (MLM) loss to learn discrete units for better contextualized representations.
In this paper, we consider how to synergistically combine the aforementioned useful factors to boost overall performance. A hierarchical information path of multilingual ASR is hypothesized and empirically examined. The proposed path is from LID prediction to acoustic Unit discovery followed by international Phonetic alphabet (IPA) sharing, then routed by MoE for final multilingual Token recognition (debudded as LUPET  [^1] LUPET in Filipino means “Awesome”.). LUPET seamlessly includes multiple pieces of information in different granularity, e.g., LID, acoustic unit, phoneme, and token. Within this path, the information of the shallow layer is beneficial for deriving those occurred in deeper layers, thus being hierarchical. Importantly, these information labels are either straightforward to obtain or derived in an SSL manner.
LUPET can be easily integrated into the vanilla ASR architecture, significantly boosting the multilingual recognition performance. We conduct extensive experiments to evaluate the proposed LUPET. Experiments on 10 languages of the Common Voice [24] show that, compared to the vanilla method, our proposed LUPET gives 19.73% and 12.27% relative averaged word error rate reductions with CTC and attention decoding, respectively. Furthermore, LUPET beats all previous baselines and exhibits superior performance on high-resource languages.
Figure 1: The overall architecture of our proposed LUPET multilingual ASR. LUPET information path unfolds with the encoder layers. { $Enc^{s}$, $Enc^{lm}$, $Enc^{um}$, $Enc^{d}$} represent shallow, lower-middle, upper-middle, deep layers, respectively. $Enc^{s}$ and $Enc^{um}$ are used for LID and IPA phoneme prediction. $Enc^{lm}$ performs acoustic unit discovery with a random-projection quantizer, where $\mathbb{C}$ denotes the codebook for vector quantization (VQ). $Enc^{d}$ denotes conformer layers modified with MoE which consists of four experts and a router. All trapezoid modules refer to linear projection.
2 LUPET Method
This section starts with a brief introduction of a vanilla E2E multilingual ASR architecture and then illustrates how to incorporate the proposed LUPET information path.
2.1 Vanilla E2E Multilingual ASR
The E2E multilingual ASR architecture we adopt is the hybrid CTC-Attention conformer [25, 26]. It consists of three components, namely encoder, decoder, and CTC [27] layer. The encoder takes an acoustic feature sequence $\mathbf{X}=\{\mathbf{x}_{t}\}_{t=1}^{T}$ as input and converts it to hidden representation $\mathbf{H}=\{\mathbf{h}_{t}\}_{t=1}^{T^{{}^{\prime}}}$, where $T$ and $T^{{}^{\prime}}$ denote the number of original frames and the sub-sampled frames. The $\mathbf{H}$ is then forwarded to two classification branches for predicting the token sequence $\mathbf{Y}=\{y_{u}\in\mathcal{V}\}_{u=1}^{U}$, where $\mathcal{V}$ is a shared multilingual vocabulary set built by BPE [28] and $U$ denotes the number of tokens. One classification branch, i.e., decoder, conditions $\mathbf{H}$ to autoregressively compute token-level posterior $p(y_{u}|\mathbf{H},y_{1:u-1})$ via a cross attention mechanism. The attention loss is given as
$\mathcal{L}_{Attn}=-\sum_{u=1}^{U}logp(y_{u}|\mathbf{H},y_{1:u-1}).\vspace{-2mm}$ (1)
Another classification branch, i.e., the CTC layer, simultaneously derives the frame-level posteriors $p(\mathbf{z}_{t}|\mathbf{h}_{t})$. The CTC loss is formulated as follows:
$\mathcal{L}_{CTC}=CTC(\mathbf{Z},\mathbf{Y})=-\sum_{Z\in B^{-1}(\mathbf{Y})}\sum_{t=1}^{T^{{}^{\prime}}}logp(\mathbf{z}_{t}|\mathbf{h}_{t}),$ (2)
where $\mathbf{z}_{t}$ is the logits over $\mathcal{V}\cup\emptyset$ and $B^{-1}$ is the inverse function that gives all the valid alignment paths between input sequence $\mathbf{H}$ and output sequence $\mathbf{Y}$. The blank token $\emptyset$ is specially designed by CTC for aligning $\mathbf{H}$ and $\mathbf{Y}$. $\mathbf{Z}=\{\mathbf{z}_{t}\}_{t=1}^{T^{{}^{\prime}}}$ represents one possible alignment path. The training objective of hybrid CTC-Attention is a linear combination of Eq. 1 and Eq. 2:
$\mathcal{L}_{CTC-Attn}=(1-\lambda)\mathcal{L}_{Attn}+\lambda\mathcal{L}_{CTC},$ (3)
where $\lambda$ is a coefficient to control the weight of CTC loss.
2.2 Incorporating LUPET
Recent studies have pointed out (1) Both LID and phoneme information are beneficial for multilingual training [11, 14]; (2) Design language-specific modules to process language-specific information is useful to reduce language interference [17, 19]; (3) The success of self-supervised cross-lingual representation learning applied in ASR [22, 29]. Although these factors' effectiveness has been verified separately, how to synergistically combine them to contribute a better solution from a unified perspective remains as an open question. The proposed LUPET provides a novel view from the multilingual hierarchical information path. From LID to acoustic unit followed by phoneme then go through MoE routing to the final token, the information that occurred in the early position of the path is assumed to contribute to the prediction of later information of the path.
As shown in Fig. 1, the full encoder is composed of { $Enc^{s}$, $Enc^{lm}$, $Enc^{um}$, $Enc^{d}$}, from shallow layers to deep layers. LUPET information path unfolds with the encoder layers. The shallow layers of encoder $Enc^{s}$ are used to identify the spoken language. Denote the output of $Enc^{s}$ as shallow representations $\mathbf{H}^{s}$. $\mathbf{H}^{s}$ is then projected to the LID logits $\mathbf{Z}^{lid}$ via a linear transformation. The logits dimension $dim(\mathbf{Z}^{lid})=$$LID+1$, where $1$ represents a special blank token for CTC as mentioned in Sec. 2.1. The LID prediction loss is formulated as:
$\mathcal{L}_{lid}=CTC(\mathbf{Z}^{lid},LID_{seq}),$ (4)
where the sequential LID labels $LID_{seq}$ are constructed by repeating the single LID label to the number of output tokens.
The predicted LID information then is propagated to subsequent layers of the encoder via self-conditioning
$\mathbf{H}^{s^{\prime}}=\mathbf{H}^{s}+LIN(\mathbf{Z}^{lid}),$ (5)
where $LIN$ denotes a linear layer to keep the hidden dimension and $\mathbf{H}^{s^{\prime}}$ is the input representations for the next encoder layer.
The lower-middle layers of encoder $Enc^{lm}$ are utilized to perform acoustic unit discovery. Similar to BEST-RQ [29], a random-projection quantizer including a projection matrix $Proj^{u}$ and codebook $\mathbb{C}$ is applied and none of the parameters are trainable. Vector quantization (VQ) is carried out on acoustic features $\mathbf{X}$ to produce discrete labels
$\mathbf{Lab}_{u}=\mathbb{C}(Proj^{c}(Sub(\mathbf{X})))),$ (6)
where $Sub$ represents the subsample operation. The output of $\mathbb{C}$ are the indices of the nearest code vectors in the codebook.
With probability $p$, $\mathbf{X}$ is randomly masked to feed the encoder. Masked language modeling (MLM) is then performed to predict $\mathbf{Lab}_{u}$. Denote $\mathbf{H}^{lm}_{M}$ as the output representation of $Enc^{lm}$, where the under-script $M$ means having Mask($\mathbf{X}$) as input. Let $\mathbf{mi}$ be the masked indices on $\mathbf{H}^{lm}_{M}$, the MLM loss can be written as:
$\mathcal{L}_{mlm}=CE(Proj^{u}(\mathbf{H}^{lm}_{M}[\mathbf{mi}]),\mathbf{Lab}_{u}[\mathbf{mi}]),$ (7)
where $Proj^{u}$ projects the hidden dimension to the size of the codebook and MLM loss is the cross-entropy ($CE$) between logits over the codebook and labels at the masked positions.
Discrete acoustic units discovered by $Enc^{lm}$ are expected to facilitate pronunciation learning to incorporate phonetic information for subsequent layers. Similar to LID prediction, the output representation $\mathbf{H}^{um}$ of upper-middle encoder $Enc^{um}$ is used to predict phoneme sequence $IPA$. $\mathbf{Z}^{ipa}$, projected by $\mathbf{H}^{um}$, is the logits over IPA phonemes and an additional blank token. Eq. 8 gives the loss of IPA prediction:
$\mathcal{L}_{ipa}=CTC(\mathbf{Z}^{ipa},IPA).$ (8)
Following Eq. 5, self-conditioning is similarly applied to obtain $\mathbf{H}^{um^{\prime}}$ to propagate the predicted phonetic information.
$\mathbf{H}^{um^{\prime}}=\mathbf{H}^{um}+LIN(\mathbf{Z}^{ipa})$ (9)
Lastly, the deep layers of encoder $Enc^{d}$ are modified with MoE. Multiple FFN experts and a routing network are included in the MoE structure. The language self-condition representation ($LIN(\mathbf{Z}^{lid})$ in Eq. 5) is regarded as the LID embedding to feed the routing network. The output of the routing network is a softmax distribution over the number of experts. Followed [17], the top-2 experts with the highest probabilities are dynamically routed to process each frame based on the frame-level LID information from shallow encoder layers.
To incorporate the hierarchical information, from LID, acoustic unit, phoneme, and token, the ultimate objective function of LUPET is given as a linear combination of Eq. (3, 4, 7, 8):
$\mathcal{L}_{LUPET}=\mathcal{L}_{CTC-Attn}+w_{1}\mathcal{L}_{lid}+w_{2}\mathcal{L}_{mlm}+w_{3}\mathcal{L}_{ipa},$ (10)
where $w_{1},w_{2},w_{3}$ are the weights of the corresponding losses.
Table 1: Training and testing hours of 10 languages from the Common Voice 13.0 corpus in our experiments.
LID en fr es zh it
Train 2279.98 872.19 448.45 359.91 286.61
Test 26.90 26.08 26.65 30.44 26.27
LID ru pt tr nl tt
Train 178.78 125.35 69.08 73.82 19.95
Test 15.73 11.91 12.05 14.58 5.70
3 Experimental Setup
3.1 Dataset
The 10 languages, namely English (en), French (fr), Spanish (es), Chinese (zh), Italian (it), Russian (ru), Portuguese (pt), Turkish (tr), Dutch (nl) and Tatar (tt) from the public available Common Voice 13.0 [24] are selected for our multilingual ASR experiments. The language coverage includes high-resource languages, e.g., English with around 2,280 hours of training data, and low-resource languages, e.g., Tatar with only about 20 hours. The detailed training and testing statistics are listed in Tab. 1. Note that zh includes Mandarin, Taiwanese, and Cantonese. A standard text normalization (the same as in Whisper [2]) is applied to all the transcriptions of the dataset.
3.2 Multilingual ASR Configurations
3.2.1 Vanilla
The vanilla model adopts a hybrid CTC-Attention architecture. The encoder has 12 conformer layers with 8 attention heads and 512 hidden dimensions, while the decoder has 6 transformer layers [30]. The CTC weight $\lambda$ in Eq. 3 is set to $0.3$. The input acoustic feature to the network is the typical 80-dimensional log-Mel filterbank. The output vocabulary used is derived from Whisper's tokenizer. This tokenizer was obtained by BPE using UTF-8 bytes of the entire training dataset of Whisper.
3.2.2 LUPET
Compared to vanilla, the encoder architecture has several modifications by incorporating LUPET. The output positions of { $Enc^{s}$, $Enc^{lm}$, $Enc^{um}$, $Enc^{d}$} are at the {3-th,6-th,9-th,12-th} layer of the original encoder, respectively. When MLM loss takes effect, acoustic feature $\mathbf{X}$ is randomly masked consecutive 20 frames with probability $p=0.01$. Codebook $\mathbb{C}$ of the random-projection quantizer has a size of 8192 and a dimension of 16. IPA sequence per-utterance is obtained using an open-sourced toolkit phonemizer  [^2] https://github.com/bootphon/phonemizer. In each layer of $Enc^{d}$, MoE including 8 FFN experts is used to replace the end-FFN of the original conformer layer. In Eq. 10, the weight coefficients $w_{1}$, $w_{2}$, and $w_{3}$ are set to 0.3, 0.07, and 0.3, respectively.
3.2.3 Baselines
Several baselines are used for comparison: (1) Mono, monolingual ASR with vanilla architecture is trained per-language. (2) Oracle_LID, append the pre-known LID embedding to input acoustic feature for multilingual ASR training. (3) MoE [17], keep the $Enc^{d}$ of LUPET and remove other auxiliary losses, hidden representation $\mathbf{H}^{um}$ is used as the input to the routing network. (4) LID_SC [11], LID prediction by CTC and LID information self-conditioning (SC) are performed. (5) Whisper, whisper-large-v2 with oracle LID is used for decoding.
3.3 Training Scheme and Evaluation Metric
We implement the vanilla and our proposed LUPET methods on the Wenet toolkit [31]. The model is trained with Adam [32] optimizer with a learning rate (LR) of $31e-$ 3. LR schedule has a warmup step of $15000$. Batch size is set to $12$ with $accum\_grad=16$. 8 V100 GPUs are used for DDP training. Each multilingual model is trained for 50 epochs and each monolingual model is trained for 100 epochs. If not specified otherwise, MLM takes effect from epoch 5 to 30. The final model for decoding is obtained by averaging the $10$ best models with the lowest validation losses. Character error rate (CER) for Chinese and word error rate (WER) for other languages are used to measure the system's performance.
4 Results and Analysis
Table 2: WER (%) results of different systems on 10 languages of Common Voice by CTC greedy decoding. In the LUPET block, the backslash /  represents the ablation study by removing the following component, where U = acoustic unit discovery, P = IPA sharing, and L = LID prediction. $w_{2}=1$ means the weight coefficient of $\mathcal{L}_{mlm}$ is set as 1. Uto50ep means U takes effect until 50 epochs.
Model en fr es zh it ru pt tr nl tt avg avg w/o tt avg 5high avg 5low
Mono 13.03 12.51 9.37 13.01 11.15 11.55 11.16 25.73 19.34 83.62 21.05 14.09 11.81 30.28
Vanilla 13.50 13.33 9.74 13.71 10.56 16.34 12.07 23.49 13.72 36.78 16.32 14.05 12.17 20.48
Oracle_LID 12.69 12.08 8.51 12.8 9.07 13.64 9.39 20.12 11.72 30.29 14.03 12.22 11.03 17.03
LID_SC 12.94 12.24 8.84 13.46 9.36 14.72 10.89 22.53 12.74 34.19 15.19 13.08 11.37 19.01
MoE 12.86 12.81 9.23 12.67 9.91 13.56 10.26 20.55 11.46 30.47 14.38 12.59 11.50 17.26
LUPET 11.75 11.79 8.22 12.41 8.81 10.95 8.95 17.71 11.32 29.12 13.10 11.32 10.60 15.61
LUPET / U 12.33 12.32 8.73 12.42 9.45 10.58 9.62 18.00 10.86 27.76 13.21 11.59 11.05 15.36
LUPET / P 12.35 12.22 8.67 12.31 9.39 11.92 10.51 21.92 12.08 27.23 13.86 12.37 10.99 16.73
LUPET / UP 12.71 12.38 8.82 12.16 9.54 11.9 10.34 20.89 11.72 26.47 13.69 12.27 11.12 16.26
LUPET / LU 11.96 12.02 8.46 12.21 9.08 10.99 9.44 19.49 11.03 31.90 13.66 11.63 10.75 16.57
LUPET $w_{2}=1$ 11.80 11.86 8.54 12.33 9.10 12.38 10.25 17.85 10.30 33.11 13.75 11.60 10.73 16.78
LUPET Uto50ep 11.72 12.09 8.40 12.73 9.02 11.90 9.98 19.06 11.96 31.35 13.82 11.87 10.79 16.85
Figure 2: Relative WER changes of different systems to monolingual systems on 10 languages by CTC greedy decoding.
4.1 Performance Comparison to Monolingual System
A desirable multilingual ASR is expected to present better performance than its monolingual counterparts (Mono). As shown in Fig. 2, LUPET and other baselines are used to compare performance to monolingual systems on the test sets of 10 languages. The x-axis follows an order from high-resource to low-resource languages. The y-axis denotes the relative WER to Mono, where the more negative value represents the lower WER. It is clear to see all curves are basically decreasing except the peak at ru. The decreasing trend is straightforward as the more low-resourced languages can achieve more significant performance gains. Multilingual training brings degradation to the Russian (ru) language in most cases. We speculate this may be due to the compromised phenomenon from Russian (ru) to Tarta (tt). The language tt achieves above 60% relative WER reduction via multilingual training at the cost of side effects on language ru, which shares the same language family as tt.
It is worth noted that the recognition performance of Vanilla system on four high-resource languages (en, fr, es, zh) cannot surpass the corresponding monolingual system, demonstrating the general compromised phenomenon towards low-resource languages in multilingual training. Oracle_LID system brings consistent improvements over Vanilla across all languages, which proves the benefits of LID information. Both MoE and LID_SC also show obviously better performance than Vanilla, while being inferior to Oracle_LID. LUPET outperforms all other baselines, serving as the only system that gives WER reduction on all languages compared to Mono. The advantage of LUPET is highlighted by the superior performance on high-resource languages, which largely mitigates the compromised phenomenon during multilingual training.
4.2 LUPET's Effectiveness Verification
Tab. 2 presents the WER results of different systems on 10 languages. Averaged WER are calculated for overall comparison. The top-5 languages, with more training data, are roughly referred to as high-resources, while the remaining five are low-resources. Having similar observation from Fig. 2, LUPET gives significantly better performance on all languages compared to other baselines. The benefits of LID prediction and MoE routing structure have been well verified by system LID_SC and MoE.
To further illustrate the effectiveness of LUPET, several ablation studies are carried out to investigate the remained components. By removing U (acoustic unit discovery) from LUPET, it can be clearly observed that WERs on high-resource languages consistently increase. Contrary to high-resource, some low-resource languages especially for tt, achieve somewhat improvements. It demonstrates the quality of discrete units that discovered by MLM is related to the amount of data. Hence, MLM can usually bring positive gains to high-resource languages. With a longer MLM effective period (Uto50ep), low-resource languages have obvious performance degradation. The gains towards high-resource gradually converge to the language en. When increasing the weight coefficient of $\mathcal{L}_{mlm}$ ($w_{2}=1$), inferior results compared to the original LUPET setting are presented for most of languages.
Disabling both U and P (IPA sharing prediction) exhibits worse results. Tab. 2 provides two comparison views to understand the independent effect of the component P. (1) LUPET / LU can be seen as ``MoE + P". Comparing it with MoE, IPA sharing is found to be beneficial for all languages except tt. (2) When only removing P from LUPET, WERs on all languages basically degrade. Low-resource languages clearly give the worse results, where tr increase the absolute WER above 4%. One possible reason is that the independent U would lead to information loss due to the masking mechanism in MLM, especially for low-resource languages. In LUPET, with the help of intermediate phoneme prediction (P), the loss of information is largely mitigated, thus not presenting much worse results on low-resource languages.
Table 3: Averaged WER (%) of different systems by attention decoding. Note that Whisper-large-v2 decodes in a greedy manner, while other systems utilize beam search with beam_size=20.
Model avg avg w/o tt avg 5high avg 5low
Whisper 20.86 11.52 13.21 28.52
Mono 17.89 10.02 9.52 26.26
Vanilla 10.43 8.78 8.88 11.99
Oracle_LID 9.20 7.89 8.31 10.08
LID_SC 10.22 8.62 8.62 11.83
MoE 9.61 8.25 8.46 10.76
LUPET 9.15 7.77 7.95 10.35
4.3 Results on Attention Decoding
Tab. 3 presents the averaged WER metrics of different systems by attention decoding. Not surprisingly, LUPET exhibits the overall best performance, especially on high-resource languages, and significantly outperforms its CTC decoding counterpart by 3.94% absolute WER. Whisper is introduced as an external reference. As can be seen, the zero-shot performance of Whisper is easily surpassed even by Mono, illustrating the importance of in-domain training. Furthermore, it is noted that the overall performance gaps between systems in attention decoding are far less than that of CTC decoding, e.g., comparing Oracle_LID and LUPET. We hypothesize that the attention decoder served as a language model may help overfit the pattern in the specific domain. This also explains why attention decoding clearly outperforms CTC decoding in our experiments.
5 Conclusions
This paper presents a novel view to seamlessly incorporate hierarchical information into multilingual ASR. Multiple information in different granularity, i.e., LID, acoustic unit, phoneme, and token, form a path LUPET that unfolds with encoder layers. Experiments carried out on 10 languages of Common Voice corpus illustrate the effectiveness of LUPET, even outperforming the system with oracle LID information. Different components in LUPET are proved to be useful in ablation studies. It is found that the acoustic unit discovery and phoneme prediction significantly help the recognition on high-resource languages, largely mitigating the compromised phenomenon.